if the distance between points p1(sinx,cosx)and p2(cosx,y)is square root of 2,what is the value of y?

Answer by stanbon(75887)

if the distance between points p1(sinx,cosx)and p2(cosx,y)is square root of 2,what is the value of y?
-------
(sin-cos)^2 + (cos-y)^2 = 2
---
sin^2-2sin*cos+cos^2 + (cos-y)^2 = 2
---------
1 - sin(2x) + (cos-y)^2 = 2
--------
(cos-y)^2 = -1-sin(2x)
----
cos-y = sqrt(-1-sin(2x))
y = cos(x)-sqrt(-1-sin(2x))
=========================
Cheers,
Stan H.
------------------